A slow-pitch softball diamond is actually a square. The distance from home plate to 1st base is 65 feet. How far is it from home plate to 2nd base? Give exact answer.
Give the approximate distance to 3 decimal places.

Answer by stanbon(75887)

A slow-pitch softball diamond is actually a square. The distance from home plate to 1st base is 65 feet. How far is it from home plate to 2nd base? Give exact answer.
Give the approximate distance to 3 decimal places.
------------------------------
Draw the picture.
The distance from home plate to 2nd base
is the hypotenuse of a right triangle
with legs = 65 ft.
-----
h^2 = 65^2 + 65^2
h^2 = 2*65^2
h = 65*sqrt(2)
h = 91.92 ft.
================
Cheers,
Stan H.